1. A normal distribution has u=80 and o=10. What is the probability of randomly selecting the following scores?

a) x > 75

b) x < 85

c) between the mean and a score of 90

d) between the mean and a score of 50

e) 75 < x < 85

Show all calculations

2. Determine the z-score value in each of the following scenarios:

a) what z-score value separates the top 8% of a normal distribution from the bottom 92%?

b) what z-score value separates the top 72% of a normal distribution from the bottom 28%?

c) what z-scores values represents the middle 90% of the values in a normal distribution?


Show all calculations


1
Expert's answer
2021-06-02T14:12:26-0400

1.

"\mu=80 \\

\sigma=10"

a)

"P(x>75) = 1 -P(x<75) \\

=1 -P(Z<\frac{75-80}{10}) \\

= 1 -P(Z< -0.5) \\

= 1 -0.3085 \\

= 0.8085"

b)

"P(x<85) = P(Z< \frac{85-80}{10}) \\

= P(Z<0.5) \\

= 0.6914"

c)

"P(80<x<90) =P(x<90) -P(x<80) \\

=P(Z< \frac{90-80}{10}) -P(Z< \frac{80-80}{10}) \\

= P(Z< 1) -P(Z<0) \\

= 0.8413 -0.5 \\

= 0.3413"

d)

"P(50<x<80) = P(x<80) -P(x<50) \\

=P(Z< \frac{80-80}{10}) -P(Z< \frac{50-80}{10}) \\

= P(Z<0) -P(Z< -3) \\

= 0.5 -0.00135 \\

= 0.49865"

e)

"P(75<X<85) = P(X<85) -P(X<75) \\

= P(Z< \frac{85-80}{10}) -P(Z< \frac{75-80}{10}) \\

=P(Z<0.5) -P(Z< -0.5) \\

= 0.6914 -0.3085=0.3829"

2.

a) P(Z<z*) = 0.92

z*=1.4051

b) P(Z<z*) = 0.28

z*=-0.5828

c)

"P(z^*_1<Z<z^*_2)=0.90 \\

0.5 - \frac{0.90}{2}=0.05 \\

P(Z<z^*_1) = 0.05 \\

z^*_1=1.65 \\
z^*_2=-1.65"